Question 2 5 pts 4.33-kg soccer ball rolling eastward at a speed of 2.74 m/s is kicked so that it reverses direction and attains a speed of 6.35 m/s. the duration of the interaction is 55.33 ms, what is the average force on the ball by the players foot, in N? A If Note: Do enter three significant figures, do not enter units and do not pay attention to the silly things Canvas is doing to your answer like adding zeros on the end
Answer #1

according to newton's second law we have

average force = dp/dt

F = m(v - u) / t

= 4.33 ( 6.35 - (-2.74)) / 0.05533

= 711.36 N

so the average force = 711.36 N